How long does it take electrons to get from a car battery to the starting motor? Assume the current is 250 A and the electrons travel through a copper wire with cross-sectional area 0.19 cm2 and length 0.79 m. The number of charge carriers per unit volume is 8.49·1028 m-3.
Answer #1

Given that

The current is (i) = 250 A

The electrons travel through a copper wire with cross-sectional area(A) = 0.19 cm2 =0.19*10-4m2

The length is (L) =0.79 m.

The number of charge carriers per unit volume is(n) = 8.49·1028 m-3.

The drift velocity is given by

vd =J/ne =i/Ane-----(1)

Now the time taken for electrons to get from a car battery to the starting motor is given by

t =L/Vd =LAne/i =(0.79m)(0.19*10-4m2)(8.49·1028 m-3)(1.6*10-19C)/250A

                        =0.000815*105s =815.583s =13.593min